Tim lives in San Francisco and runs a business that sells pianos. In an average year, he receives $851,000 from selling pianos. Of this sales revenue, he must pay the manufacturer a wholesale cost of $476,000; he also pays wages and utility bills totaling $281,000. He owns his showroom; if he chooses to rent it out, he wil receive $71,000 in rent per year. Assume that the value of this showroom does not depreciate over the year. Also, if Tim does not operate this piano business, he can work as an accountant, receive an annual salary of $34,000 with no additional monetary costs, and rent out his showroom at the $71,000 per year rate. No other costs are incurred in running this piano business Identify each of Tims costs in the following table as either an implicit cost or an explicit cost of selling pianos. Implicit Cost Explicit Cost The wholesale cost for the pianos that Tim pays the manufacturer The salary Tim could earn if he worked as an accountant The wages and utility bills that Tim pays The rental income Tim could receive if he chose to rent out his showroom Complete the following table by determining Tims accounting and economic profit of his piano business Profit (Dollars) Accounting Profit Economic Profit
Answer #1

a. The wholesale cost for the pianos that Tim pays the manufacturer - Explicit cost

b. The salary Tim could earn if he worked as an accountant - Implicit cost

c. The wages and utility bills that Tim pays - Explicit cost

d. The rental income Tim could receive if he chose to rent out his showroom - Implicit cost

(Explicit costs are direct expenses by the firm whereas Implicit cost are the opportunity cost due to lose of other alternative)

Accounting profit = Total revenue - explicit cost = ($851,000) - ($476,000 + $281,000) = 851,000 - 757,000 = $94,000

Economic profit = Accounting profit - implicit cost = $94,000 - ($71,000 + $34,000) = 94,000 - 105,000 = -$11,000